OpenStudy (anonymous):

What value of x makes the equation true? –4.9 = –5.7 + x A. x = –10.6 B. x = –0.8 C. x = 0.8 D. x = 10.6

OpenStudy (jacobciezki):

There's two ways to do this, depending on which math you are taking. The first or 'elementary school' way would be to plug in all of the choices. There's also the 'algebra' way, which would be to find out what x is without using the choices. -4.9 = -5.7 + x Add* (not subtract) the 5.7 to both sides 0.8 = x And there you have it! The answer is C, but now you know how to find it when you don't have choices. *If you were wondering why you add 5.7, it's because it cancels out the -5.7 - one of the most common mistakes is subtracting it, but you see why that's wrong, right?

OpenStudy (anonymous):

-4.9+5,7=x; 0.8=x (C)

OpenStudy (anonymous):

yea

OpenStudy (jacobciezki):

so its c